40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 630     Entry time: Thu Jul 3 13:12:32 2008
Author: Rob, Yoichi, John 
Type: Update 
Category: Locking 
Subject: More oscillations 
Bounce/ roll filters were added to the short degrees of freedom to reduce the effect of the 24Hz line seen on Tuesday night.

However last night saw the arrival of a new oscillation at ~34Hz. This may be the second harmonic of the MOS roll mode. Reducing the arm offset would cause this oscillation to ring up and break the lock (first plot). This effect was repeatable.

No signal was seen in the oplevs or osems which leads us to rule out alignment problems, at least for now.

Although one can clearly see DARM_ERR increasing as arm power increases adding a resonant gain in the DARM loop had no effect.

We also noticed that x arm transmission was significantly more noisy than the Y (second plot). And showed greater coherence with the increase in DARM noise. Investigations showed that the PD was not the source of the difference.

Turning up the MC gain seems to help a little.

We're now looking at POX as a candidate for RF_CARM (third plot).
Attachment 1: LOL.png  46 kB  Uploaded Thu Jul 3 15:29:05 2008  | Hide | Hide all
LOL.png
Attachment 2: NoisyX080702.png  14 kB  Uploaded Thu Jul 3 15:31:57 2008  | Hide | Hide all
NoisyX080702.png
Attachment 3: POXforCARM080702.png  15 kB  Uploaded Thu Jul 3 15:32:15 2008  | Hide | Hide all
POXforCARM080702.png
ELOG V3.1.3-